OpenStudy (anonymous):

Use scientific notation to divide the following two numbers. Express the answer using scientific notation; retain at least three decimal places. (7.56*10^4)/(6.92*10^2)

OpenStudy (anonymous):

\[\frac{7.56*10^4}{6.92*10^2}=\frac{7.56}{6.92} \times \frac{10^4}{10^2}=1.09 \times 10^2\]

OpenStudy (anonymous):

is that the whole problem

OpenStudy (mertsj):

Yes it is. Except you were supposed to have three decimal places so it would be 1.092 x10^2